OpenStudy (anonymous):

4x^2+1>-4x I don't know how to solve the first interval

jimthompson5910 (jim_thompson5910):

4x^2+1>-4x 4x^2+1+4x>-4x+4x 4x^2+4x+1 > 0 now you need to solve 4x^2+4x+1 = 0 for x

OpenStudy (imstuck):

Bring everything onto one side of the inequality sign, then set it equal to 0. factor it and you're good!

OpenStudy (anonymous):

okay I got (2x+1)(2x+1) but I'm guessing thats not right?

jimthompson5910 (jim_thompson5910):

you should have (2x+1)(2x+1) = 0 so x = ??

OpenStudy (anonymous):

-1/2

jimthompson5910 (jim_thompson5910):

notice how (2x+1)(2x+1) is the same as (2x+1)^2

jimthompson5910 (jim_thompson5910):

(2x+1)^2 is either positive or 0 it will never be negative

jimthompson5910 (jim_thompson5910):

so (2x+1)^2 > 0 is true for all real numbers x

jimthompson5910 (jim_thompson5910):

well actually excluding x = -1/2

OpenStudy (anonymous):

I'm confused on what the answer is

jimthompson5910 (jim_thompson5910):

The answer is \[\Large \left(-\infty, -\frac{1}{2}\right)\cup\left(-\frac{1}{2},\infty\right)\] because we're just excluding the value -1/2 from the solution set

OpenStudy (anonymous):

it says its wrong, but i noticed i wrote the original equation wrong its a \[\ge \] instead of >

OpenStudy (anonymous):

how does that change it?

jimthompson5910 (jim_thompson5910):

oh, then the solution set is all real numbers

jimthompson5910 (jim_thompson5910):

in interval notation that would be \[\Large \left(-\infty,\infty\right)\]

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome